# Face-Human-Bench: A Comprehensive Benchmark of Face and Human Understanding for Multi-modal Assistants

### v2

@@ B.2.1 Prompt Templates for Different Settings
Zero-Shot (ZS) The prompt template used for the zero-shot setting is shown in Table 14.
Hints (H) The prompt template for experiments with hints is shown in Table 15.
@@
Hints and Vanilla CoT Instructions (H+VCoT) The prompt template for experiments with hints and vanilla CoT instructions is shown in Table 16.
@@
Hints and Task-Specific Instructions With One-Stage Framework (H+1TCoT) The prompt template for one-stage experiments with hints and task-specific CoT instructions is shown in Table 17.
@@
Hints and Task-Specific Instructions With Two-Stage Framework (H+2TCoT) The prompt template for two-stage experiments with hints and task-specific CoT instructions is shown in Table 18.
@@ B.2.2 Prompt Used for Choice Extraction
The prompt used for choice extraction is shown in Table 19.
@@ B.2.3 Hints and Task-specific CoT Instructions
Hints and task-specific CoT instructions for each L3 ability are shown in Table LABEL:tab:long-table.
@@
Age / Please (1) analyze the facial age characteristics of the person in the image and (2) provide a possible age number that you think is appropriate. Note: Please do not respond with ”I can’t determine the exact age”; just provide the number you think is closest.
@@
B.3 More Details on the Experiments for RQ2
@@ B.3.1 Unexplored L3 Abilities
Face/Human Attribute Recognition These two tasks include a large number of binary classification labels (40 labels in CelebA for face and 14 labels in WIDER Attribute for human). Using evaluation protocols designed for specialist models to fully assess the performance of MLLMs would result in huge computational costs. Additionally, many attribute labels have ambiguous semantics that are difficult to define accurately, such as ”attractive,” ”big lips,” and ”big nose.”
@@
Social Relationship Recognition Existing evaluation datasets for specialist models, such as PIPA and PISC, are not suitable for directly evaluating the social relationship recognition abilities of MLLMs. Because many annotations in these datasets are ambiguous (it is often impossible to assert that people in an image belong to one relationship category rather than another) and semantically overlapping (for example, “couple” is a separate category, but a married couple is semantically also part of “family”). Fine-tuned specialist models can still learn to classify under unclear standards, but for zero-shot learning MLLMs, the lack of clear definitions greatly reduces performance. It is worth noting that in Face-Human-Bench, when we use PISC to construct problems for evaluating the social relationship recognition ability of MLLMs, we manually review and remove problems with ambiguous options or potentially non-unique answers, allowing us to assess accurately.
@@ B.3.2 Explored L3 Abilities
We provide the prompt templates for directly evaluating L3 abilities on public datasets from the face and human community in Tables 21, 22, 23, 24, 25, 26, 27, 28 and 29. During testing, the options will be randomly shuffled.
@@
Are the people in the two photos the same person? A. Yes B. No Answer with the option’s letter from the given choices directly, and don’t provide other additional explanations. Only one option is correct.
@@ C.1 Face-Human-Bench (English)
We provide the visualization of the L2 and L3 results in Figures 8, 9 and 10.
@@ C.2 Face-Human-Bench (Chinese)
Table 30 shows the performance of all evaluated MLLMs at different levels of abilities on the Human-Face-Bench (Chinese). We further compare the performance of different MLLMs on English and Chinese versions of the Face-Human-Bench, as shown in Figure 11. Models are sorted with the ascending order of average performance.
Table 30: Zero-shot scores of MLLMs on the hierarchical Face-Human-Bench (CN). The highest scores for open-source and closed-source MLLMs are marked in blue and green respectively.
@@
Model Attr. Action RPU CC Mean SRR IR Mean Re-ID Face Human Per. Rea. Overall
@@ C.3 Correlation Between Abilities
The correlation coefficient matrix for L3 is shown in Figure 12. Pay particular attention to the ability correlations highlighted in the red boxes.
@@ C.4 Relative Position of Targets
Table 31 presents the performance differences of MLLMs across different relative positions of targets, under the three face understanding abilities and human attribute recognition.
@@ C.5 CoT prompting
Based on Table 32, we explore the main reasons for the performance improvements of GPT-4o in each ability at L3, as shown in Figure 13.
@@
InternVL -Chat-v1.2-Plus H+2TCoT 88.0 62.3 72.0 54.0 63.0 58.0 66.5 62.3 94.0 66.0 56.0 78.0 56.0 70.0
@@
Model Setting Attr. Action RPU CC Mean SRR IR Mean Re-ID Face Human Per. Rea. Overall
@@
InternVL -Chat-v1.2-Plus H+2TCoT 87.0 92.0 58.0 51.3 54.6 72.0 92.0 82.0 80.0 69.1 79.1 75.8 71.8 74.1
@@
Abilities with performance improvements mainly due to vanilla CoT instructions include facial attribute recognition, deepfake detection, face anti-spoofing, occluded face recognition, human attribute recognition, action recognition, relative position understanding, and social relationship recognition. Comparison of outputs from H and H + VCoT settings is shown in Tables 33, 34, 35 and 36.
Abilities with performance improvements mainly due to 1-stage task-specific CoT instructions include compound expression recognition and crowd Counting. Comparison of outputs from H + VCoT and H + 1TCoT settings is shown in Table 37.
Abilities with performance improvements mainly due to 2-stage task-specific CoT instructions include compound expression recognition and cross-pose face recognition. Comparison of outputs from H + 1TCoT and H + 2TCoT settings is shown in Table 38.
@@ C.6 Specialist Models Significantly Outperforming MLLMs
We list the early specialist models used for comparison in Table 39.
@@
ACC 44.55 DLP-CNN (Li et al., 2017b)
Deepfake FF++ ACC 82.01 XceptionNet Chollet (2017)
Spoofing SiW-Mv2 ACER 9.40 SRENet Guo et al. (2022)
Basic FR LFW ACC 99.50 R50 (He et al., 2016) + CosFace (Wang et al., 2018) + CASIA-WebFace (Yi et al., 2014)
@@
Action HICO-DET mAP 19.81 ConsNet (Liu et al., 2020)
Counting ShTech-A MAE 110.20 MCNN (Zhang et al., 2016)
Re-ID Market1501 ACC 95.26 LightMBN (Herzog et al., 2021)
@@ Appendix D Potential Bias for Demographic Characteristics
Do MLLMs contain potential biases? Specifically, do their performances vary based on the demographic characteristics of the input faces? Existing works, such as constructing the RFW (Wang et al., 2019) and BFW (Robinson et al., 2020) datasets, have explored racial biases in face recognition systems. Inspired by these works, we investigate whether MLLMs exhibit different face recognition abilities across different racial groups.
We transform face pairs from the Caucasian, African, Asian, and Indian subsets of the RFW dataset into face recognition problems similar to those in Face-Human-Bench. The test results of the three best-performing open-source models in our main experiments are presented in Table 40, revealing the racial bias of MLLMs in face recognition ability. The performance of Caucasians is the best for each model, significantly surpassing that of other racial groups. In our future work, we will systematically evaluate the performance variations of MLLMs on samples with different demographic characteristics.
@@
Appendix E Privacy Protection
Face-Human-Bench can also be used to evaluate privacy protection. In some scenarios, we want MLLMs to refuse to answer certain questions related to faces and humans. In such cases, lower performance on the Face-Human-Bench indicates a higher success rate in privacy protection on this information. Table 41 presents a comparison of the performance between APIs provided by OpenAI and Azure OpenAI. Note that Azure OpenAI primarily offers security and enterprise-grade services. GPT-4V and GPT-4o from Azure OpenAI show significant performance degradation in age estimation and expression recognition. Here are some example outputs:
  • “I cannot determine the age of the person in the photo with the information provided.”
  • “I’m sorry, but the image is too blurry to make an accurate assessment of the person’s age.”
  • “I don’t have enough visual information from the image provided to accurately determine the emotion being expressed by the person.”
  • “I’m unable to determine the person’s expression due to the blurred face. Based on the available data, I cannot select a correct answer from the provided options.”
@@ Appendix E Privacy Protection
Model Attr. Action RPU CC Mean SRR IR Mean Re-ID Face Human Per. Rea. Overall
@@ Appendix F A demonstration of How to Enhance Multi-Modal Assistant Responses with Specialist Models
In Figure 14, we use media forensics as an application scenario to demonstrate how specialist models can improve the response quality of a multimodal assistant. Path 1 directly uses the MLLM to generate responses, while Path 2 introduces a well-trained specialist model for deepfake detection to determine whether there are digital artifacts on the faces in the image. By using the output of the specialist model to enhance the prompt, Path 2 ultimately allows the MLLM to provide more accurate responses.

### v3

@@ B.2.1 Prompt Templates for Different Settings
Zero-Shot (ZS) The prompt template used for the zero-shot setting is shown in Table˜15.
Hints (H) The prompt template for experiments with hints is shown in Table˜16.
@@
Hints and Vanilla CoT Instructions (H+VCoT) The prompt template for experiments with hints and vanilla CoT instructions is shown in Table˜17.
@@
Hints and Task-Specific Instructions With One-Stage Framework (H+1TCoT) The prompt template for one-stage experiments with hints and task-specific CoT instructions is shown in Table˜18.
@@
Hints and Task-Specific Instructions With Two-Stage Framework (H+2TCoT) The prompt template for two-stage experiments with hints and task-specific CoT instructions is shown in Table˜19.
@@ B.2.2 Prompt Used for Choice Extraction
The prompt used for choice extraction is shown in Table˜20.
@@ B.2.3 Hints and Task-specific CoT Instructions
Hints and task-specific CoT instructions for each L3 ability are shown in LABEL:tab:long-table.
@@
Age / Please (1) analyze the facial age characteristics of the person in the image and (2) provide a possible age number that you think is appropriate. Note: Please do not respond with "I can’t determine the exact age"; just provide the number you think is closest.
@@
B.3 More Details on the Experiments for Q2
@@ B.3.1 Unexplored L3 Abilities
Face/Human Attribute Recognition These two tasks include a large number of binary classification labels (40 labels in CelebA for face and 14 labels in WIDER Attribute for human). Using evaluation protocols designed for specialist models to fully assess the performance of MLLMs would result in huge computational costs. Additionally, many attribute labels have ambiguous semantics that are difficult to define accurately, such as "attractive," "big lips," and "big nose."
@@
Social Relationship Recognition Existing evaluation datasets for specialist models, such as PIPA and PISC, are not suitable for directly evaluating the social relationship recognition abilities of MLLMs. Because many annotations in these datasets are ambiguous (it is often impossible to assert that people in an image belong to one relationship category rather than another) and semantically overlapping (for example, “couple" is a separate category, but a married couple is semantically also part of “family"). Fine-tuned specialist models can still learn to classify under unclear standards, but for zero-shot learning MLLMs, the lack of clear definitions greatly reduces performance. It is worth noting that in Face-Human-Bench, when we use PISC to construct problems for evaluating the social relationship recognition ability of MLLMs, we manually review and remove problems with ambiguous options or potentially non-unique answers, allowing us to assess accurately.
@@ B.3.2 Explored L3 Abilities
We provide the prompt templates for directly evaluating L3 abilities on public datasets from the face and human community in Tables˜22, 23, 24, 25, 26, 27, 28, 29 and 30. During testing, the options will be randomly shuffled.
@@ C.1 Face-Human-Bench (English)
We provide the visualization of the L2 and L3 results in Figures˜8, 9 and 10.
@@ C.2 Face-Human-Bench (Chinese)
Table˜31 shows the performance of all evaluated MLLMs at different levels of abilities on the Human-Face-Bench (Chinese). We further compare the performance of different MLLMs on English and Chinese versions of the Face-Human-Bench, as shown in Figure˜11. Models are sorted with the ascending order of average performance.
Table 31: Zero-shot scores of MLLMs on the hierarchical Face-Human-Bench (CN). The highest scores for open-source and closed-source MLLMs are marked in blue and green respectively. The scores in the “random" row are theoretical values.
@@
Model Attr. Action RPU $C$ Mean SRR IR Mean Re-ID Face Human Per. Rea. Overall
@@ C.3 Correlation Between Abilities
The correlation coefficient matrix for L3 is shown in Figure˜12. Pay particular attention to the ability correlations highlighted in the red boxes.
@@ C.4 Relative Position of Targets
Table˜32 presents the performance differences of MLLMs across different relative positions of targets, under the three face understanding abilities and human attribute recognition.
@@ C.5 CoT prompting
Based on Table˜33, we explore the main reasons for the performance improvements of GPT-4o in each ability at L3, as shown in Figure˜13.
@@
InternVL
-Chat-v1.2-Plus
H+2TCoT 88.0 62.3 72.0 54.0 63.0 58.0 66.5 62.3 94.0 66.0 56.0 78.0 56.0 70.0
@@
Model Setting Attr. Action RPU $C$ Mean SRR IR Mean Re-ID Face Human Per. Rea. Overall
@@
InternVL
-Chat-v1.2-Plus
H+2TCoT 87.0 92.0 58.0 51.3 54.6 72.0 92.0 82.0 80.0 69.1 79.1 75.8 71.8 74.1
@@
Abilities with performance improvements mainly due to vanilla CoT instructions include facial attribute recognition, deepfake detection, face anti-spoofing, occluded face recognition, human attribute recognition, action recognition, relative position understanding, and social relationship recognition. Comparison of outputs from H and H + VCoT settings is shown in Tables˜34, 35, 36 and 37.
Abilities with performance improvements mainly due to 1-stage task-specific CoT instructions include compound expression recognition and crowd Counting. Comparison of outputs from H + VCoT and H + 1TCoT settings is shown in Table˜38.
Abilities with performance improvements mainly due to 2-stage task-specific CoT instructions include compound expression recognition and cross-pose face recognition. Comparison of outputs from H + 1TCoT and H + 2TCoT settings is shown in Table˜39.
@@ C.6 Specialist Models Significantly Outperforming MLLMs
We list the early specialist models used for comparison in Table˜40.
@@
ACC 44.55 DLP-CNN [29]
Deepfake FF++ ACC 82.01 XceptionNet [8]
Spoofing SiW-Mv2 ACER 9.40 SRENet [17]
Basic FR LFW ACC 99.50
R50 [21]
+ CosFace [58]
+ CASIA-WebFace [66]
@@
Action HICO-DET mAP 19.81 ConsNet [39]
Counting ShTech-A MAE 110.20 MCNN [71]
Re-ID Market1501 ACC 95.26 LightMBN [22]
C.7 Statistical Significance for Face-Human-Bench
The core contribution of this paper is the introduction of Face-Human-Bench, a benchmark designed to evaluate the performance of MLLMs. We aim to confirm that the proposed benchmark can significantly reflect performance differences across models.
Since the overall score is computed as a weighted average from several subset accuracies, it is crucial to validate the effectiveness of tests for each individual subset.
For each subset $j$, we posit the null hypothesis $H_{0}^{(j)}$: All models exhibit identical true performance on subset $j$, with the observed variations attributed solely to measurement noise.
First, we calculate the unbiased sample variance across 25 models:
$S_{j}^{2}=\frac{1}{24}\sum_{i=1}^{25}\left(\text{Acc}_{ij}-\bar{\text{Acc}}_{j}\right)^{2},$ (1)
where $\bar{\text{Acc}}_{j}$ represents the mean accuracy for subset $j$. Next, we compute the expected variance under $H_{0}^{(j)}$:
$\sigma_{j}^{2}=\frac{\bar{\text{Acc}}_{j}(1-\bar{\text{Acc}}_{j})}{K_{j}}.$ (2)
Here, $K_{j}$ is the size of subset $j$. We derive the chi-squared statistic that quantifies the discrepancy between the observed and expected variances:
$\chi_{j}^{2}=\frac{24\cdot S_{j}^{2}}{\sigma_{j}^{2}}.$ (3)
Under $H_{0}^{(j)}$, $\chi_{j}^{2}$ follows a chi-squared distribution with 24 degrees of freedom. Across all subsets, we observe that $\chi_{j}^{2}\gg 24$ and the p-values are close to 0, indicating rejection of $H_{0}^{(j)}$. This demonstrates that our benchmark can significantly reflect the differences in model performance.
@@ Appendix D Potential Bias for Demographic Characteristics
Do MLLMs contain potential biases? Specifically, do their performances vary based on the demographic characteristics of the input faces? Existing works, such as constructing the RFW [59] and BFW [50] datasets, have explored racial biases in face recognition systems. Inspired by these works, we investigate whether MLLMs exhibit different face recognition abilities across different racial groups.
We transform face pairs from the Caucasian, African, Asian, and Indian subsets of the RFW dataset into face recognition problems similar to those in Face-Human-Bench. The test results of the three best-performing open-source models in our main experiments are presented in Table˜41, revealing the racial bias of MLLMs in face recognition ability. The performance of Caucasians is the best for each model, significantly surpassing that of other racial groups. In our future work, we will systematically evaluate the performance variations of MLLMs on samples with different demographic characteristics.
@@
Appendix E Societal Impacts
@@
Model Attr. Action RPU $C$ Mean SRR IR Mean Re-ID Face Human Per. Rea. Overall
@@
Our work aims to inspire the community to build multi-modal assistants with improved response quality and broadened application scope by providing a comprehensive and scientific evaluation of MLLMs’ face and human understanding abilities. This reflects the positive societal impacts of our research.
At the same time, we recognize that in some instances, personal privacy must be adequately protected. Therefore, it is important to proactively limit MLLMs’ ability to understand facial and bodily information to prevent the improper extraction of private information.
Our proposed Face-Human-Bench can also be used to evaluate privacy protection. In such scenarios, we want MLLMs to refuse to answer certain questions related to faces and humans. In such cases, lower performance on the Face-Human-Bench indicates a higher success rate in privacy protection on this information. Table˜42 presents a comparison of the performance between APIs provided by OpenAI and Azure OpenAI. Note that Azure OpenAI primarily offers security and enterprise-grade services. GPT-4V and GPT-4o from Azure OpenAI show significant performance degradation in age estimation and expression recognition. Here are some example outputs:
  • “I cannot determine the age of the person in the photo with the information provided."
  • “I’m sorry, but the image is too blurry to make an accurate assessment of the person’s age."
  • “I don’t have enough visual information from the image provided to accurately determine the emotion being expressed by the person."
  • “I’m unable to determine the person’s expression due to the blurred face. Based on the available data, I cannot select a correct answer from the provided options."
@@ Appendix F A demonstration of How to Enhance Multi-Modal Assistant Responses with Specialist Models
In Figure˜14, we use media forensics as an application scenario to demonstrate how specialist models can improve the response quality of a multimodal assistant. Path 1 directly uses the MLLM to generate responses, while Path 2 introduces a well-trained specialist model for deepfake detection to determine whether there are digital artifacts on the faces in the image. By using the output of the specialist model to enhance the prompt, Path 2 ultimately allows the MLLM to provide more accurate responses.